2. Assume that the pdf of the random variable x is uniform in the interval (10, 12) and y = x^3.

(a) Find fy (y).

(b) Find E{y}.
Answer #1

a)

here CDF of Y: F(y)=P(Y<y)=P(X3 <y)=P(X<y1/3)=\int_{10}^{y^{1/3}}(f(x) dx=\int_{10}^{y^{1/3}}(1/2) dx=(x/2)|y1/310 =(1/2)*(y1/3-101/3)

therefore pdf f(y)=(d/dy)*F(y)=(1/6)*y-2/3 for 1000<y<1728

b)

E(Y)=E(X3)=\int_{10}^{{12} x3(f(x) dx =\int_{10}^{{12}x3/2 dx =x4/8 |1210 =1342